When released from rest, a 200 g block slides down the path shown below, reaching the bottom with a speed of 4.2 m/s. How much work does the force of friction do (in J)? Distance of hill is 1.9m
Answer #1

Force due to gravity, Fg = 0.2 x 9.8 = 1.96 N

Work, Wg = 1.96 x 1.9 = 3.72 J

Change in kinetic energy, dE = 0.5 x 0.2 x 4.2^2 = 1.76 J

Work due to friction, Wf = 1.76 - 3.72 = - 1.96 J

(In case you want only magnitude then avoid negative sign while entering)

Comment in case any doubt please rate my answer...